# Neuronal correlations shape the scaling behavior of memory capacity and nonlinear computational capability of recurrent neural networks

### v2

@@ Abstract
Reservoir computing is a powerful framework for real-time information processing, characterized by its high computational ability and quick learning, with applications ranging from machine learning to biological systems. In this paper, we demonstrate that the memory capacity of a reservoir recurrent neural network scales sublinearly with the number of readout neurons. To elucidate this phenomenon, we develop a theoretical framework for analytically deriving memory capacity, attributing the decaying growth of memory capacity to neuronal correlations. In addition, numerical simulations reveal that once memory capacity becomes sublinear, increasing the number of readout neurons successively enables nonlinear processing at progressively higher polynomial orders. Furthermore, our theoretical framework suggests that neuronal correlations govern not only memory capacity but also the sequential growth of nonlinear computational capabilities. Our findings establish a foundation for designing scalable and cost-effective reservoir computing, providing novel insights into the interplay among neuronal correlations, linear memory, and nonlinear processing.
@@ I Introduction
Over the past decade, deep learning has advanced significantly, with generative AI in particular beginning to transform many aspects of daily life. Although deep neural networks possess the capacity to represent highly complex functions, they typically require substantial computational resources, large volumes of training data, and long training times. These high computational demands pose a major obstacle to applying deep learning in real-time computing on edge-computing devices [1].
@@
Reservoir computing (RC) is a machine learning framework for efficiently training large scale recurrent neural networks (RNNs), referred to as ”reservoirs” in the context of RC [2, 3]. In contrast to conventional training methods such as backpropagation through time (BPTT), RC optimizes only readout weights and leaves the remaining weights fixed, which enables quick and low-cost learning. In spite of its training simplicity, RC has high computational performance and a broad range of applications for processing time-series data [4]. RC is not restricted to RNNs, and a wide variety of dynamical systems can be utilized as reservoirs under appropriate conditions. In particular, RC using a real physical system such as soft matter and optical systems, so called ”physical reservoir computing”, has been intensively studied in recent years [5].
@@
In a typical RC setting, the readout connections linking reservoir units to output units are sparse [5]. This means that the size of the reservoir ($N$) is substantially larger than the number of readout units ($L$) that connect to the output units. Empirically, the performance of RC improves as an increase in $L$. However, implementing a large number of readout connections can be resource-intensive, especially in physical RC implementations. Consequently, understanding the relationship between $L$ and the computational ability of a reservoir is essential for constructing cost-effective RC. In particular, performance in RC depends on both the reservoir’s memory capacity [6] and nonlinear processing capability [7], and the number of readout units must be carefully chosen to balance these two aspects according to the demands of a given task. Despite its practical importance, a practical principle for determining an appropriate $L$ remains largely unexplored. In this study, we aim to address this gap by elucidating how memory capacity and nonlinear computational ability scale with $L$, with the particular focus on the role of neuronal correlations.
@@
We demonstrate, for the first time, that the memory capacity of an RNN increases monotonically with $L$, although its growth rate progressively declines (Fig.1). This sublinear scaling of memory capacity cannot be explained by the previous theoretical works assuming a scaling regime where $L$ is of order one relative to $N$, i.e. $L\sim O(1)$ [8, 9], because the memory capacity in this regime is proportional to $L$, as will be shown later. To bridge the gap between the observation and theory, we develop a novel theory for analytically deriving memory capacity for $L\sim O(\sqrt{N})$. Using our theoretical framework, we demonstrate that the correlations between reservoir neurons, despite their very weak magnitude scaling as $O(1/\sqrt{N})$, play an important role in the declining growth rate of memory capacity. Furthermore, numerical simulations show that the ability to perform more complex nonlinear tasks emerges supralinearly and sequentially as $L$ increases. These findings indicate that the number of readout connections should be tailored to the specific memory and nonlinearity requirements of the task at hand.
@@ II Model setup
We investigate computational capacity of RC using a large random RNN as a reservoir (Fig.1(a)), which is known as Echo state network (ESN) [2], one of the canonical model for RC  [10, 11]. The state of $i$-th neuron at discrete time $t$ is described by the variable, $x_{i}(t)$. The activation function $\phi$ is assumed to be an odd saturating sigmoid function satisfying $\phi^{\prime}(0)=1$, $\phi^{\prime}(x)>0$, $\phi^{\prime\prime}(x)\leq 0\ (x\geq 0)$ and $\phi(\pm\infty)=\pm 1$. The time evolution of a reservoir RNN is determined by the difference equation,
@@
The macroscopic dynamical behavior of the large random RNN described by Eq.(1) has been extensively studied by means of dynamical mean-field theory (DMFT) [12]. In the absence of the inputs and noise, i.e. $\sigma_{s}^{2}=\sigma_{n}^{2}=0$, the RNN exhibits phase transition from zero fixed-point to chaotic dynamics at $g=1$ in the large network size limit [12, 13]. In the presence of inputs ($\sigma_{s}^{2}>0$ or $\sigma_{n}^{2}>0$), chaotic variability is typically suppressed [8, 13, 14, 15, 9, 16], although it has recently been shown that in RNNs incorporating gating mechanisms similar to those in gated recurrent units (GRUs), external inputs can instead induce chaos [17].
The output of the reservoir RNN is defined as $z(t)=\sum_{i=1}^{L}w_{i}x_{i}(t)$, where $L$ is the number of readout neurons. We assume that the readout neurons are sparse, i.e. $L\ll N$. According to the RC framework, the output weights, $\{w_{i}\}_{i=1}^{L}$, are optimized to minimize the time-averaged squared error between the output signals, $\{z(t)\}_{t=1}^{T}$, and the desired signals, $\{f(t)\}_{t=1}^{T}$.
@@ III Memory capacity
and the angular bracket denotes time averaging. A high value of $M_{d}$ indicates that the reservoir can accurately output its input signal from $d$ steps prior. It is shown that $0\leq M_{d}\leq 1$ holds true [6, 7]. Memory capacity is defined as the infinite sum of $M_{d}$, but it has been proven to be finite and bounded by the number of readout neurons, i.e. $0\leq MC\leq L$  [6, 7]. In particular, for RC with a linear activation function operating in noise-free conditions ($\sigma_{n}^{2}=0$), memory capacity is equivalent to $L$ independently of $g$ and $\sigma_{s}^{2}$  [6].
Fig.1(b) illustrates the relationship between $L$ and the memory capacity for the reservoir RNN described by Eq.(1), obtained through numerical simulations. As shown, memory capacity exhibits monotonic growth, but its growth rate gradually diminishes as $L$ increases. Consequently, memory capacity can be characterized as a sublinear function of $L$.
Figure 1: (a) Overview of RC utilizing a random RNN receiving input signals and inherent neuronal noise. The shaded region represents the readout neurons. (b) Numerical simulations for memory capacity of the reservoir RNNs with network size $N=1000$, noise intensity $\sigma_{n}^{2}=0.1^{2}$, and activation function $\phi(x)=\tanh(x)$. Shaded area represents mean±$\pm$±std of direct numerical simulations for 10 different network and input signal realizations. The sum of $M_{d}$ is calculated up to $d=1000$. The dashed line indicates the theoretical upper bound of memory capacity, $MC=L$. Simulation time length is $T=10^{4}$.
@@ IV Theoretical analysis of memory capacity
where both $\alpha$ and $\tilde{\sigma}_{s}^{2}$ are $O(1)$. The parameter $\alpha$ represents the relative magnitude of $L$ compared to $\sqrt{N}$. In contrast to previous studies [18, 8, 9], where $L\sim O(1)$, our model employs a significantly larger number of readout neurons, although the output connections remain sparse, as $L=\alpha\sqrt{N}\ll N$
By simple calculation, $M_{d}$, defined in Eq.(2), can be expressed as
@@
where the elements of $\bm{a}_{d}\in\mathbb{R}^{L}$ and $C\in\mathbb{R}^{L\times L}$ are respectively $(\bm{a}_{d})_{i}\equiv\langle s(t-d)x_{i}(t)\rangle$ and $C_{ij}\equiv\langle x_{i}(t)x_{j}(t)\rangle$ ($i,j$ indicate the indices of the readout neurons) [7]. The inverse of the covariance matrix, $C^{-1}$, prevents us from making progress on analytical calculation. Previous theoretical studies circumvented this issue by assuming the off-diagonal entries of $C$ are zero, that is, ignoring neuronal correlations [18, 8, 9]. However, this approximation results in linear scaling of memory capacity, $MC\propto L$, conflicting with our numerical simulations as shown in Fig.1(b).
We noticed that from the scale assumption Eq.(3), the diagonal entries of $C$ are $O(1)$, whereas the non-diagonal ones are $O(1/\sqrt{N})$ (See Appendix.A.2). Hence, the non-diagonal elements are much smaller than the diagonal ones, which motivates us to perform Neumann series expansion,
$C^{-1}=(D+\tilde{C})^{-1}=\sum_{n=0}^{\infty}D^{-1}\left(-\tilde{C}D^{-1}\right)^{n},$ (5)
where $D_{ij}\equiv\delta_{ij}\langle x_{i}(t)^{2}\rangle$ and $\tilde{C}_{ij}\equiv(1-\delta_{ij})\langle x_{i}(t)x_{j}(t)\rangle$. This approach allows us to circumvent the issue of the inverse matrix and enables an analytical calculation of $M_{d}$.
In the large network size limit, we can assume self-averaging for $M_{d}$, $\lim_{N\to\infty}M_{d}=\lim_{N\to\infty}[M_{d}]$, where the square bracket denotes the average over network realizations, known as quenched average [19]. Consequently, substituting Eq.(5) into Eq.(4) and taking quenched average of $M_{d}$, we analytically obtain memory capacity:
@@
where the value of $[\langle x_{i}^{2}\rangle]$ can be obtained by solving a self-consistent equation,
@@
In deriving this result, it is necessary to calculate quenched averages of quantities involving neuronal correlations, such as
@@
However, these quantities cannot be evaluated using DMFT because it reduces the original $N$-body system to an effective single-body description, thereby neglecting inter-neuronal correlations. To overcome this limitation, we adopt the dynamical cavity method, a technique originally introduced by Clark et.al [20] from statistical physics to the analysis of neural networks. A detailed description of this theory is provided in Appendix.A.1 and A.2, followed by the full derivation of Eq.(6) in Appendix.B.1. In the following, we use the shorthand notation, $\langle f(x)\rangle_{*}\equiv\langle f(x)\rangle_{x\sim\mathcal{N}(0,[\langle x_{i}^{2}\rangle])}$.
[…]
@@
where $[\langle x_{i}^{2}\rangle]$ is determined by solving
@@
We have confirmed that the analytical values of $MC$ are consistent with the numerical simulations, exhibiting a sublinear function of $L$, for $L$ below a threshold (Fig.2). When $L$ exceeds the threshold, the analytical values rapidly deviate from the simulated ones and eventually diverge (dashed lines in Fig.2). This is because Neumann series expansion, Eq.(5), does not converge as the spectrum norm $\|\tilde{C}D^{-1}\|$ exceeds one for large $L$. To evaluate the upper bound of $L$ ensuring convergence of the series expansion, it is enough to evaluate $\|\tilde{C}D^{-1}\|$, but this is a challenging task. Instead, we derive the sufficient condition for the expansion series to converge (see Appendix.B.1), and the corresponding upper bounds of $L$ satisfying this condition are indicated by star marks in Fig.2.
@@ V Neuronal correlations shape the sublinear scaling of memory capacity
Having obtained the analytical solution of memory capacity, we now investigate the underlying reason why memory capacity grows sublinearly with respect to $L$. To quantify the extent to which the growth rate of memory capacity decays, we define a decay rate,
@@
By this definition, if $MC(L)$ scales linearly with $L$, then $r(L)=1$ holds for all $L$. In contrast, $r(L)$ decays from unity when $MC(L)$ is a sublinear function of $L$. As depicted in Fig.2, $r(L)$ decays for various parameter combinations, confirming the sublinear scaling of memory capacity with respect to $L$.
Employing the solution of memory capacity (Eq.(6)), we obtain an analytical form of the decay rate (see Appendix.B.2 for derivation):
@@
×$\hskip 60.0pt\times\frac{1-(g\langle\phi^{\prime}(x)\rangle_{*})^{2}}{1-(g\langle\phi^{\prime}(x)\rangle_{*})^{2n+2}}.$

### v3

# Neuronal correlations shape the scaling behavior of memory capacity and nonlinear computational capability of reservoir recurrent neural networks

@@ Abstract
Reservoir computing is a powerful framework for real-time information processing, characterized by its high computational ability and quick learning, with applications ranging from machine learning to biological systems. In this paper, we investigate how the computational ability of reservoir recurrent neural networks (RNNs) scales with an increasing number of readout neurons. First, we demonstrate that the memory capacity of a reservoir RNN scales sublinearly with the number of readout neurons. To elucidate this observation, we develop a theoretical framework for analytically deriving memory capacity that incorporates the effect of neuronal correlations, which have been ignored in prior theoretical work for analytical simplicity. Our theory successfully relates the sublinear scaling of memory capacity to the strength of neuronal correlations. Furthermore, we show this principle holds across diverse types of RNNs, even those beyond the direct applicability of our theory. Next, we numerically investigate the scaling behavior of nonlinear computational ability, which, alongside memory capacity, is crucial for overall computational performance. Our numerical simulations reveal that as memory capacity growth becomes sublinear, increasing the number of readout neurons successively enables nonlinear processing at progressively higher polynomial orders. Our theoretical framework suggests that neuronal correlations govern not only memory capacity but also the sequential growth of nonlinear computational capabilities. Our findings establish a foundation for designing scalable and cost-effective reservoir computing, providing novel insights into the interplay among neuronal correlations, linear memory, and nonlinear processing.
@@ I Introduction
Over the past decade, deep learning has advanced significantly [1], with generative AI in particular beginning to transform many aspects of daily life [2]. Although deep neural networks possess the capacity to represent highly complex functions [3, 4], they typically require substantial computational resources, large volumes of training data, and long training times [5]. These high computational demands pose a major obstacle to applying deep learning in real-time computing on edge-computing devices [6].
@@
Reservoir computing (RC) is a machine learning framework for efficiently training large-scale recurrent neural networks (RNNs), referred to as ”reservoirs” in the context of RC [7, 8]. In contrast to conventional training methods such as backpropagation through time (BPTT), RC optimizes only readout weights and leaves the remaining weights fixed, which enables quick and low-cost learning. Despite its training simplicity, RC has high computational performance and a broad range of applications for processing time-series data [9]. RC is not restricted to RNNs, and a wide variety of dynamical systems can be utilized as reservoirs under appropriate conditions. In particular, RC using a real physical system such as soft matter and optical systems, so called ”physical reservoir computing”, has been intensively studied in recent years [10].
@@
In a typical RC setting, the readout connections linking reservoir units to output units are sparse [10]. This means that the size of the reservoir ($N$) is substantially larger than the number of readout units ($L$) that connect to the output units. The performance of RC empirically improves as $L$ increases because this increases the expressive power of the reservoir output [11]. In contrast, implementing a large number of readout connections can be resource-intensive, especially in physical RC implementations. Consequently, understanding the relationship between $L$ and the computational ability of a reservoir is essential for constructing cost-effective RC. Despite its practical importance, however, a guiding principle for determining an appropriate $L$ has yet to be fully explored.
@@
In this study, we aim to address this gap by elucidating how memory capacity [12] and nonlinear computational ability [13] – two key determinants of a reservoir’s computational power – scale with $L$, with the particular focus on the role of neuronal correlations. We demonstrate, for the first time, that the memory capacity of an RNN increases monotonically as increasing $L$, with its growth rate progressively declining (Fig.1). This sublinear scaling of memory capacity cannot be explained by the previous theoretical works assuming a scaling regime where $L$ does not scale with $N$ , i.e. $L\sim O(1)$ [14, 15], because the memory capacity in this regime exhibits linear scaling, as will be shown later. To bridge the gap between our observations and existing theory, we develop a novel theory for analytically deriving memory capacity for $L\sim O(\sqrt{N})$. Using our theoretical framework, we demonstrate that the correlations between reservoir neurons, despite being very weak with a magnitude scaling as $O(1/\sqrt{N})$, play an important role in the declining growth rate of memory capacity. Furthermore, numerical simulations show that the ability to perform more complex nonlinear tasks emerges supralinearly and sequentially as $L$ increases. These findings indicate that the number of readout connections should be tailored based on the specific memory and nonlinearity requirements of the task at hand.
@@ II Model setup
We investigate computational capacity of RC using a large random RNN as a reservoir (Fig.1(a)), which is known as Echo State Network (ESN) [7], one of the canonical models for RC  [16, 17]. The state of $i$-th neuron at discrete time $t$ is described by the variable $x_{i}(t)$. The activation function $\phi$ is assumed to be an odd saturating sigmoid function satisfying $\phi^{\prime}(0)=1$, $\phi^{\prime}(x)>0$, $\phi^{\prime\prime}(x)\leq 0\ (x\geq 0)$ and $\phi(\pm\infty)=\pm 1$. The time evolution of a reservoir RNN is determined by the difference equation,
@@
The macroscopic dynamical behavior of the large random RNN described by Eq.(1) has been extensively studied by means of dynamical mean-field theory (DMFT) [18]. In the absence of the inputs and noise, i.e. $\sigma_{s}^{2}=\sigma_{n}^{2}=0$, the RNN exhibits phase transition from zero fixed-point to chaotic dynamics at $g=1$ in the large network size limit $N\to\infty$ [18, 19]. In the presence of inputs ($\sigma_{s}^{2}>0$ or $\sigma_{n}^{2}>0$), chaotic variability is typically suppressed [14, 19, 20, 21, 15, 22], while it has recently been shown that in RNNs incorporating gating mechanisms similar to those in gated recurrent units (GRUs), external inputs can instead induce chaos [23].
The output of the reservoir RNN is defined as $z(t)=\sum_{i=1}^{L}w_{i}x_{i}(t)$, where $L$ is the number of readout neurons. We assume that the readout neurons are sparse, i.e., $L\ll N$. According to the RC framework, the output weights, $\{w_{i}\}_{i=1}^{L}$, are optimized to minimize the time-averaged squared error between the output signals, $\{z(t)\}_{t=1}^{T}$, and the desired signals, $\{f(t)\}_{t=1}^{T}$. Since this is a least-squares problem, the solution for $\bm{w}\in\mathbb{R}^{L}$ is explicitly given by
$\bm{w}=(XX^{\top})^{-1}X\bm{f},$ (2)
where the matrix $X\in\mathbb{R}^{L\times T}$ contains the history of readout neuron activities throughout simulation time length $T$, and the vector $\bm{f}\in\mathbb{R}^{T}$ denotes the target time-series.
@@ III Memory capacity
and the angular bracket denotes time averaging. Here, the optimized readout weights $\bm{w}$ are obtained by substituting f(t)=s(td)f(t)=s(t-d) into Eq.(2). A high value of $M_{d}$ indicates that the reservoir can accurately output its input signal from $d$ steps prior. It is shown that $0\leq M_{d}\leq 1$ holds true [12, 13]. Memory capacity is defined as the infinite sum of $M_{d}$, but it has been proven to be finite and bounded by the number of readout neurons, i.e. $0\leq MC\leq L$  [12, 13]. In particular, for RC with a linear activation function operating in noise-free conditions ($\sigma_{n}^{2}=0$), memory capacity is equivalent to $L$, independently of $g$ and $\sigma_{s}^{2}$  [12].
Fig.1(b) illustrates the relationship between $L$ and the memory capacity for the reservoir RNN described by Eq.(1), obtained through numerical simulations. As shown, memory capacity grows monotonically, but its growth rate gradually diminishes as $L$ increases. Consequently, memory capacity can be characterized as a sublinear function of $L$.
[…]
@@ IV Theoretical analysis of memory capacity
where both $\alpha$ and $\tilde{\sigma}_{s}^{2}$ are of $O(1)$. The parameter $\alpha$ represents the relative magnitude of $L$ compared to $\sqrt{N}$. In contrast to previous studies [24, 14, 15], where $L\sim O(1)$, our model employs a significantly larger number of readout neurons, although the output connections remain sparse, as $L=\alpha\sqrt{N}\ll N$. It should be noted that while our theoretical analysis in this study is conducted under this assumption, our theoretical framework can be extended to more general case where $L\sim O(N^{c})$ and $\sigma_{s}^{2}\sim O(1/N^{c})$ for $0\leq c\leq 1/2$ (see Appendix B.2).
Substituting the optimal solution for $\bm{w}$ from Eq.(2) into the definition of $M_{d}$ (Eq.(3)) yields:
@@
where the elements of $\bm{a}_{d}\in\mathbb{R}^{L}$ and $C\in\mathbb{R}^{L\times L}$ are respectively $(\bm{a}_{d})_{i}\equiv\langle s(t-d)x_{i}(t)\rangle$ and $C_{ij}\equiv\langle x_{i}(t)x_{j}(t)\rangle$ ($i,j$ indicate the indices of the readout neurons) [13]. The inverse of the covariance matrix, $C^{-1}$, prevents us from making progress on analytical calculation. Previous theoretical studies circumvented this issue by assuming the off-diagonal entries of $C$ are zero, that is, ignoring neuronal correlations [24, 14, 15]. However, as discussed later, this approximation results in linear scaling of memory capacity, $MC\propto L$, conflicting with our numerical simulations as shown in Fig.1(b).
We notice that from the scale assumption Eq.(4), the diagonal entries of $C$ are of $O(1)$, whereas the non-diagonal ones are of $O(1/\sqrt{N})$ (See Appendix A.2). Hence, the non-diagonal elements are much smaller than the diagonal ones, which motivates us to perform Neumann series expansion,
$C^{-1}=(C_{\rm diag}+C_{\rm nondiag})^{-1}=\sum_{n=0}^{\infty}C_{\rm diag}^{-1}\left(-C_{\rm nondiag}C_{\rm diag}^{-1}\right)^{n},$
where $(C_{\rm diag})_{ij}\equiv\delta_{ij}\langle x_{i}(t)^{2}\rangle$ and $(C_{\rm nondiag})_{ij}\equiv(1-\delta_{ij})\langle x_{i}(t)x_{j}(t)\rangle$. This approach allows us to circumvent the issue of the inverse matrix and enables an analytical calculation of $M_{d}$.
In the large network size limit, we can assume self-averaging for $M_{d}$, i.e., $\lim_{N\to\infty}M_{d}=\lim_{N\to\infty}[M_{d}]$, where the square bracket denotes the average over network realizations, known as quenched average [25]. Consequently, substituting Eq.(IV) into Eq.(5) and taking quenched average of $M_{d}$, we analytically obtain memory capacity as:
@@
where the value of $\langle x_{i}^{2}\rangle]$ can be obtained by solving a self-consistent equation (dynamical mean-field equation [18]),
@@
The memory capacity of a linear RNN with neuronal noise can be derived in a similar fashion, and this derivation also reveals a sublinear scaling (see Appendix B.4).
The derivation of these results requires calculating quenched averages of terms involving neuronal correlations, such as
@@
However, these quantities cannot be evaluated using DMFT because it reduces the original $N$-body system to an effective single-body description, thereby neglecting inter-neuronal correlations. To overcome this limitation, we adopt the dynamical cavity method, a technique originally introduced by Clark et al. [26] from statistical physics to the analysis of neural networks. A detailed description of this theory is provided in Appendix A.1 and A.2, followed by the full derivation of Eq.(7) in Appendix B.1. In the following, we use the shorthand notation, $\langle f(x)\rangle_{*}\equiv\langle f(x)\rangle_{x\sim\mathcal{N}(0,[\langle x_{i}^{2}\rangle])}$.
When the activation function is an error function, $\phi(x)=\int_{0}^{x}e^{-\frac{\pi}{4}t^{2}}dt$, we can analytically integrate both $\langle\phi^{\prime}(x)\rangle_{*}$ and $\langle\phi(x)^{2}\rangle_{*}$, yielding
@@
where $\langle x_{i}^{2}\rangle]$ is determined by solving
@@
We confirm that the analytical values of $MC$ are consistent with the numerical simulations, exhibiting a sublinear function of $L$, for $L$ below a threshold, as illustrated in Fig.2. When $L$ exceeds the threshold, the analytical values rapidly deviate from the simulated ones and eventually diverge (dashed lines in Fig.2). This divergence occurs because the Neumann series expansion (Eq.(IV)) fails to converge when the spectral norm $\|C_{\rm nondiag}C_{\rm diag}^{-1}\|$ exceeds one for large $L$. Evaluating this spectral norm to determine the convergence threshold for $L$ is a challenging task. Instead, we note that the spectral norm is bounded by the Frobenius norm $\|\cdot\|_{F}$ and that $\|C_{\rm nondiag}C_{\rm diag}^{-1}\|_{F}$ can be obtained analytically. We therefore derive a sufficient condition for the series expansion to converge (see Appendix B.1). The corresponding upper bounds of $L$ satisfying this condition are indicated by star marks in Fig.2.
@@ V Neuronal correlations shape the sublinear scaling of memory capacity
Based on our analytical solution for memory capacity, we can now investigate the underlying reason for its sublinear scaling with $L$. To quantify the extent to which the growth rate of memory capacity decays, we define a decay rate,
@@
By this definition, if $MC(L)$ scales linearly with $L$, then $r(L)=1$ holds for all $L$. In contrast, $r(L)$ decays from unity when $MC(L)$ is a sublinear function of $L$. As depicted in the inset of Fig.2, $r(L)$ consistently decays for various parameter combinations, confirming the sublinear scaling of memory capacity with respect to $L$.
From our analytical solution for memory capacity (Eq.(7)), we obtain the analytical form for the decay rate (see Appendix B.3 for the derivation) as:
@@
$\hskip 60.0pt\times\frac{1-(g\langle\phi^{\prime}(x)\rangle_{*})^{2}}{1-(g\langle\phi^{\prime}(x)\rangle_{*})^{2n+2}}.$